# Exploring the Impact of Large Language Models on Recommender Systems: An Extensive Review

### 02 Mar 2024 v2

@@ 1. Introduction
Recommendation systems are crucial for personalized content discovery, and the integration of LLMs in Natural Language Processing (NLP) is revolutionizing these systems (Huang et al., 2023a). LLMs, with their language comprehension capabilities, recommend items based on contextual understanding, eliminating the need for explicit behavioral data. The current research landscape is witnessing a surge in efforts to leverage LLMs for refining recommender systems, transforming recommendation tasks into exercises in language understanding and generation. These models excel in contextual understanding, adapting to zero and few-shot domains (Cui et al., 2022), streamlining processes, and reducing environmental impact. This paper delves into how LLMs enhance transparency and interactivity in recommender systems, continuously refining performance through user feedback. Despite challenges, researchers propose approaches to effectively leverage LLMs, aiming to bridge the gap between strengths and challenges for improved system performance. Essentially, this paper makes three significant contributions to the realm of LLMs in recommenders:
@@
POSO (Du et al., 2023), fill=bg29, text width = 22em] ] [Off-the-shelf Recommender System
@@
ChatQA (Liu et al., 2024a),fill=bg8,text width = 22em ] ] [Presonalized Recommender Systems
@@
ControlRec (Qiu et al., 2023), fill=bg5,text width = 22em ] ] [Knowledge Graph Recommender System
@@
LLMRG (Wang et al., 2019),fill=bg121,text width = 22em ] ] [ Reranking in Recommender System
@@
Ranking GPT (Zhang et al., 2024), fill = bg25,text width = 22em ] ] [ Prompts Engineered in Recommender System
@@
RecSys (Fan et al., 2023), fill = bg113,text width = 22em ] ] [ FineTuned LLMs for Recommender System
@@
INTERS (Zhu et al., 2024), fill = bg10,text width = 22em ] ] [ Evaluating LLM in Recommender System
@@ 2. Background and Related Work
In this segment, we provide a concise overview of pertinent literature concerning recommender systems and methods involving LLMs.
2.1. Recommender System
Traditional recommender systems follow Candidate Generation, Retrieval, and Ranking phases. However, the advent of LLMs like GPT  (Radford et al., 2018) and BERT  (Devlin et al., 2019) brings a new perspective. Unlike conventional models, LLMs, such as GPT and BERT, do not require separate embeddings for each user/item interaction. Instead, they use task-specific prompts encompassing user data, item information, and previous preferences. This adaptability allows LLMs to generate recommendations directly, dynamically adapting to various contexts without explicit embeddings. While departing from traditional models, this unified approach retains the capacity for personalized and contextually-aware recommendations, offering a more cohesive and adaptable alternative to segmented retrieval and ranking structures.
$2.$ Large Language Models (LLMs)
LLMs, such as BERT, GPT, Text-To-Text Transfer Transformer (T5) (Raffel et al., 2023), and Llama, are versatile in NLP. BERT is an encoder-only model with bidirectional attention, GPT employs a transformer decoder for one-directional processing, and T5 transforms NLP problems into text generation tasks. Recent LLMs like GPT-3 (Brown et al., 2020), Language Model for Dialogue Applications(LaMDA) (Thoppilan et al., 2022), Pathways Language Model (PaLM) (Chowdhery et al., 2022), and Vicuna excel in understanding human-like textual knowledge, employing In-Context Learning (ICL) (Dong et al., 2023) for context-based responses.
@@ 3.1. LLM-Powered Recommender Systems Across Domains
3.2. Off-the-shelf LLM based Recommender System
In this section we examine recommender systems that operate without tuning, specifically focusing on whether adjustments have been made to the LM.
@@ 3.2. Off-the-shelf LLM based Recommender System
$3.$ LLMs in Sequential Recommender Systems
@@ 3.5. LLMs in Personalized Recommenders Systems
Health-LLM: Jin et al. (Jin et al., 2024) introduce a framework integrating LLM and medical expertise for enhanced disease prediction and health management using patient health reports. It involves extracting informative features, assigning weighted scores by medical professionals, and training a classifier for personalized predictions. Health-LLM offers detailed modeling, individual risk assessments, and semi-automated feature engineering, providing professional and tailored intelligent healthcare.
@@
GIRL:Zheng et al. (Zheng et al., 2023) introduce GIRL, a job recommendation approach inspired by LLMs. It uses Supervised Fine-Tuning for generating Job Descriptions (JDs) from CVs, incorporating a reward model for CV-JD matching. Reinforcement Learning fine-tunes the generator with Proximal Policy Optimization, creating a candidate-set-free, job seeker-centric model. Experiments on a real-world dataset demonstrate significant effectiveness, signaling a paradigm shift in personalized job recommendation.
@@
3.6. LLMs for Knowledge Graph Enhancement in Recommender Systems
@@
LLMRG: Wang et al. (Wang et al., 2024) introduce LLM Reasoning Graphs (LLMRG) (Wang et al., 2024), utilizing LLMs to create personalized reasoning graphs for robust and interpretable recommender systems. It connects user profiles and behavioral sequences, incorporating chained graph reasoning, divergent extension, self-verification, and knowledge base self-improvement. LLMRG enhances recommendations without extra user or item information, offering adaptive reasoning for a comprehensive understanding of user preferences and behaviors.
@@ 4. Prompt Engineered LLMs in Recommender Systems
RecSys: Fan et al. (Fan et al., 2023) introduce the innovative use of prompting in tailoring LLMs for specific downstream tasks, focusing on recommendation systems (RecSys). This method involves using task-specific prompts to guide LLMs, aligning downstream tasks with language generation during pre-training. The study discusses techniques like ICL and CoT in the context of recommendation tasks within RecSys. It also covers prompt tuning and instruction tuning, incorporating prompt tokens into LLMs and updating them based on task-specific recommendation datasets.
@@ 5. Fine-tuned LLMs in Recommender Systems
Flan-T5: Kang et al. (Kang et al., 2023) from Google Brain team conducted a study to assess various LLMs with parameter sizes ranging from 250 million to 540 billion on tasks related to predicting user ratings. The evaluation encompassed zero-shot, few-shot, and fine-tuning scenarios. For fine-tuning experiments, the researchers utilized Flan-T5-Base (250 million parameters) and Flan-U-PaLM (540 billion parameters). In zero-shot and few-shot experiments, GPT-3 models from OpenAI and the 540 billion parameters Flan-U-PaLM were employed. The experiments revealed that the zero-shot performance of LLMs significantly lags behind traditional recommender models.
@@
6. Evaluating LLM in Recommender System
@@
FaiRLLM: Zhang et al. (Zhang et al., 2023a) proposes that there is a recognition of the need to assess RecLLM’s fairness. The evaluation focuses on various user-side sensitive attributes, introducing a novel benchmark called Fairness of Recommendation via LLM (FaiRLLM). This benchmark includes well-designed metrics and a dataset with eight sensitive attributes across music and movie recommendation scenarios. Using FaiRLLM, an evaluation of ChatGPT reveals persistent unfairness to specific sensitive attributes during recommendation generation.
@@
7. Challenges and Solutions
This section explores how LLMs overcome challenges in recommender systems, addressing scalability issues and emphasizing effective re-ranking strategies. The research, exemplified by Ren et al. (Ren et al., 2023), introduce the Representation Learning with LLMs for Recommendation (RLMRec) framework. RLMRec integrates representation learning with LLMs to bridge ID-based recommenders, maintaining precision and efficiency. It leverages LLMs’ text comprehension for understanding semantic elements in user behaviors and preferences, transforming textual cues into meaningful representations. The framework introduces a user/item profiling paradigm reinforced by LLMs and aligns the semantic space of LLMs with collaborative relational signals through a cross-view alignment framework.

### 19 Mar 2024 v3

@@ 1. Introduction
Recommendation systems are crucial for personalized content discovery, and the integration of Large Language Models (LLMs) in Natural Language Processing (NLP) is revolutionizing these systems (Huang et al., 2023a). LLMs, with their language comprehension capabilities, recommend items based on contextual understanding, eliminating the need for explicit behavioral data. The current research landscape is witnessing a surge in efforts to leverage LLMs for refining recommender systems, transforming recommendation tasks into exercises in language understanding and generation. These models excel in contextual understanding, adapting to zero and few-shot domains (Cui et al., 2022), streamlining processes, and reducing environmental impact. This paper delves into how LLMs enhance transparency and interactivity in recommender systems, continuously refining performance through user feedback. Despite challenges, researchers propose approaches to effectively leverage LLMs, aiming to bridge the gap between strengths and challenges for improved system performance. Essentially, this paper makes three significant contributions to the realm of LLMs in recommenders:
@@
POSO (Du et al., 2023), fill=bg29, text width = 22em] ] [Off-the-shelf Recommender Systems
@@
ChatQA (Liu et al., 2024a),fill=bg8,text width = 22em ] ] [Personalized Recommender Systems
@@
ControlRec (Qiu et al., 2023), fill=bg5,text width = 22em ] ] [Knowledge Graph-enhanced Recommender Systems
@@
LLMRG (Wang et al., 2019),fill=bg121,text width = 22em ] ] [ Reranking in Recommender Systems
@@
Ranking GPT (Zhang et al., 2024), fill = bg25,text width = 22em ] ] [ Prompt Engineered LLMs in Recommender Systems
@@
RecSys (Fan et al., 2023), fill = bg113,text width = 22em ] ] [ Fine-tuned LLMs for Recommender Systems
@@
INTERS (Zhu et al., 2024), fill = bg10,text width = 22em ] ] [ Evaluating LLMs in Recommender Systems
@@ 2. Background and Related Work
In this section, we provide a concise overview of pertinent literature concerning recommender systems and methods involving LLMs.
2.1. Recommender Systems
Traditional recommender systems follow Candidate Generation, Retrieval, and Ranking phases. However, the advent of LLMs brings a new perspective. Unlike conventional models, LLMs do not require separate embeddings for each user/item interaction. Instead, they use task-specific prompts encompassing user data, item information, and previous preferences. This adaptability allows LLMs to generate recommendations directly, dynamically adapting to various contexts without explicit embeddings. While departing from traditional models, this unified approach retains the capacity for personalized and contextually-aware recommendations, offering a more cohesive and adaptable alternative to segmented retrieval and ranking structures.
2.2. Large Language Models (LLMs)
LLMs, such as Llama (Touvron et al., 2023), GPT  (Radford et al., 2018), T5 (Raffel et al., 2023), etc., are versatile in NLP. BERT is an encoder-only model with bidirectional attention, GPT employs a transformer decoder for one-directional processing, and T5 transforms NLP problems into text generation tasks. Recent LLMs like GPT-3 (Brown et al., 2020), Language Model for Dialogue Applications (LaMDA) (Thoppilan et al., 2022), Pathways Language Model (PaLM) (Chowdhery et al., 2022), and Vicuna excel in understanding human-like textual knowledge, employing In-Context Learning (ICL) (Dong et al., 2023) for context-based responses.
@@ 3.1. LLM-Powered Recommender Systems Across Domains
3.2. Off-the-shelf LLM-based Recommender Systems
In this section we examine recommender systems that operate without tuning, specifically focusing on whether adjustments have been made to the LLM.
@@ 3.2. Off-the-shelf LLM-based Recommender Systems
3.3. LLMs in Sequential Recommender Systems
@@ 3.5. LLMs in Personalized Recommenders Systems
Health-LLM: Jin et al. (Jin et al., 2024) introduce a framework integrating LLM and medical expertise for enhanced disease prediction and health management using patient health reports. Health-LLM involves extracting informative features, assigning weighted scores by medical professionals, and training a classifier for personalized predictions. Health-LLM offers detailed modeling, individual risk assessments, and semi-automated feature engineering, providing professional and tailored intelligent healthcare.
@@
GIRL: Zheng et al. (Zheng et al., 2023) introduce GIRL, a job recommendation approach inspired by LLMs. It uses Supervised Fine-Tuning (SFT) for generating Job Descriptions (JDs) from CVs, incorporating a reward model for CV-JD matching. Reinforcement Learning fine-tunes the generator with Proximal Policy Optimization, creating a candidate-set-free, job seeker-centric model. Experiments on a real-world dataset demonstrate significant effectiveness, signaling a paradigm shift in personalized job recommendation.
@@
3.6. LLMs for Knowledge Graph-enhanced Recommender Systems
@@
LLMRG: LLM Reasoning Graphs (LLMRG) by Wang et al. (Wang et al., 2024) utilize LLMs to create personalized reasoning graphs for robust and interpretable recommender systems. LLMRG connects user profiles and behavioral sequences, incorporating chained graph reasoning, divergent extension, self-verification, and self-improvement of knowledge bases. LLMRG enhances recommendations without extra user or item information, offering adaptive reasoning for a comprehensive understanding of user preferences and behaviors.
@@ 4. Prompt Engineered LLMs in Recommender Systems
Recommender Systems in the Era of LLMs: Fan et al. (Fan et al., 2023) survey the innovative use of prompting in tailoring LLMs for specific downstream tasks, focusing on recommendation systems. They review methods involving task-specific prompts to guide LLMs, aligning downstream tasks with language generation during pre-training. The study discusses techniques like ICL and CoT in the context of recommendation tasks within RecSys. It also covers prompt tuning and instruction tuning, incorporating prompt tokens into LLMs and updating them based on task-specific recommendation datasets.
@@ 5. Fine-tuned LLMs in Recommender Systems
Flan-T5: Kang et al. (Kang et al., 2023) conduct a study to assess various LLMs with parameter sizes ranging from 250 million to 540 billion on tasks related to predicting user ratings. The evaluation encompassed zero-shot, few-shot, and fine-tuning scenarios. For fine-tuning experiments, the researchers utilized Flan-T5-Base (250 million parameters) and Flan-U-PaLM (540 billion parameters). In zero-shot and few-shot experiments, GPT-3 models from OpenAI and the 540 billion parameters Flan-U-PaLM were employed. The experiments revealed that the zero-shot performance of LLMs significantly lags behind traditional recommender models.
@@
6. Evaluating LLMs in Recommender System
@@
FaiRLLM: Zhang et al. (Zhang et al., 2023a) posit that there is a need to assess RecLLM’s fairness. The evaluation focuses on various user-side sensitive attributes, introducing a novel benchmark called Fairness of Recommendation via LLM (FaiRLLM). This benchmark includes well-designed metrics and a dataset with eight sensitive attributes across music and movie recommendation scenarios. Using FaiRLLM, an evaluation of ChatGPT reveals persistent unfairness to specific sensitive attributes during recommendation generation.